# Zero-Shot Image Segmentation via Recursive Normalized Cut on Diffusion Features

### v1

@@ 3.2 Recursive Feature Clustering
Deriving from our observation on the patch-level alignment of our chosen diffusion features, we assume that the normalized cut algorithm would provide sharp segments, each corresponding to a precise semantic concept as distinct objects would manifest as weakly connected components in a patch similarity matrix. Following this intuition, we construct an affinity matrix $\bm{W}$, by computing the cosine similarity between each pair of patches.  In addition, since NCut criterion assesses both the overall dissimilarity among different segments and the similarity within each segment, we opt to emphasize inter-segments dissimilarity by exponentiating each element with a positive integer value $\alpha$:
@@ 3.3 High-Resolution Concept Assignment
    Concept Assignment. A naive approach to obtain segmentation maps at the original image resolution consists in performing a nearest-neighbor upsampling. Despite its straightforwardness, this approach results in a blocky output structure as all pixels within the same feature patch are assigned to the same concept. Alternatively, we opt to first bilinearly upsample our low-resolution feature map $\hat{z}$ to match the original image spatial size and then proceed with the pixel/concept assignment. Specifically, for each concept $c_{i}\in\mathcal{C}$, we compute its cosine similarity with the upsampled features $\hat{z}_{up}$. This yields a similarity matrix of size $(H\times W\times K)$ where $K=|\mathcal{C}|$. Then, the assignment process simply consists in taking the argmax across the $K$ channels of the similarity matrix. The obtained segmentation map $\mathcal{S}$ is eventually refined with a pixel-adaptive refinement module [69].
@@ 4.1 Results on Zero-shot Segmentation
Model LD AX UA VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K Extra-Training    IIC [19] 9.8 - - 6.7 6.4 -    MDC [76] - - - 9.8 7.1 -    PiCIE [21] - - - 13.8 12.3 -    PiCIE+H [21] - - - 14.4 -    STEGO [20] - - - 28.2 21.0 -    ACSeg [18] 53.9 - - 28.1 - - Training-free    ReCO [53] 25.1 19.9 15.7 26.3 19.3 11.2    MaskCLIP [52] 38.8 23.6 20.6 19.6 10.0 9.8    DiffSeg [1] - - - 43.6 21.2 -    DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7    DiffCut (Ours) 65.2 56.5 34.1 49.1 30.6 44.3
With our set of default hyperparameters, DiffCut significantly outperforms all other baselines despite not relying on language dependency, auxiliary images or unsupervised adaptation. In average, our method gains +8.5 mIoU over the second best baseline. In particular, it significantly surpasses the previous state-of-the-art method in unsupervised segmentation, DiffSeg,  by +5.5 and +9.4 mIoU on COCO-Stuff and Cityscapes respectively. The effectiveness of our method is further corroborated with our qualitative results shown in Fig. 1. In comparison to DiffSeg, DiffCut provides finely detailed segmentation maps that more closely align with semantic concepts. Additional examples can be found in Supplementary J.
We note here that, as the granularity of annotations varies across target datasets, our fixed set of hyperparameters can not be in the optimal regime on each of them. Therefore, relaxing the condition on prior knowledge about the target dataset, we report in Supplementary E results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images from the target training split.
@@ 4.2 Semantic Coherence in Vision Encoders
As in [79], we collect patch representations from various vision encoders and store their corresponding target classes using the segmentation labels. Given $\hat{z}_{i}=\mathcal{E}(\mathbf{x_{1}})_{i}\in\mathbb{R}^{D_{v}}$ and $\hat{z}_{j}=\mathcal{E}(\mathbf{x_{2}})_{j}\in\mathbb{R}^{D_{v}}$, the patch representations of images $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ at respectively index $i$ and $j$, we compute their cosine similarity  and use this score as a binary classifier to predict if the two patches belong to the same class. Given $l(\mathbf{x_{1}})_{i}$ and $l(\mathbf{x_{2}})_{j}$, the labels associated to the patches, if $l(\mathbf{x_{1}})_{i,j}=l(\mathbf{x_{2}})_{p,q}$, the target value for binary classification is 1, else 0. We present in Fig. 3 the ROC curve and AUC score for our candidate models. We observe that SSD-1B UNet encoder [35] demonstrates a greater patch-level alignment than any other candidate model with an AUC score of 0.83, even surpassing DINOv2 [5]. We further exhibit the outstanding alignment between patch representations associated to semantically similar concepts with qualitative results in Fig. 4. We provide additional qualitative examples patch-level alignment in Supplementary I.
@@ 4.3 Ablation study
Table 2: Ablation Study. DiffCut recursive partitioning algorithm yields superior results than self-attention merging of DiffSeg and Automated Spectral Clustering.
@@
DiffCut vs Automated Spectral Clustering.
@@ 4.4 Model Analysis
Figure 5: Robustness of DiffCut. As $\alpha$ increases, DiffCut shows competitive results for a broad range of $\tau$ values.
Segmentation Robustness.
@@
Model LD VOC Context COCO-Object Extra-Training    ReCO [53] 25.1 19.9 15.7    ViL-Seg [82] 37.3 18.9 -    MaskCLIP [52] 38.8 23.6 20.6    TCL [50] 55.0 30.4 31.6    CLIPpy [48] 52.2 - 32.0    GroupVIT [49] 52.3 22.4 24.3    ViewCo [51] 52.4 23.0 23.5    SegCLIP [54] 52.6 24.7 26.5    OVSegmentor [83] 53.8 20.4 25.1 Training-free    CLIP-DIY [84] 59.9 19.7 31.0    FreeSeg-Diff [85] 53.3 - 31.0    DiffCut 63.0 24.6 36.0
@@ Appendix A Recursive Normalized Cut on Diffusion Features
  • Encode the image $\mathcal{I}$ with the VQ-encoder of the diffusion model: $z=\mathcal{E}_{\text{VAE}}(\mathcal{I})$
@@ Appendix B Impact of PAMR
Appendix C Mask Upsampling
@@
Appendix D Visual Encoders KMeans Comparison
@@
Appendix E Hyperparameter $\tau$ tuning
@@
Appendix F Image Noising
@@
Appendix G DiffCut Robustness
@@
Appendix H Visualization of the effect of $\tau$
@@
Appendix I Semantic Coherence in Vision Encoders
@@
Appendix J Additional Visualization
@@
Appendix K Datasets Licenses

### v2

# DiffCut: Catalyzing Zero-Shot Semantic Segmentation with Diffusion Features and Recursive Normalized Cut

@@ 3.2 Recursive Feature Clustering
Deriving from our observation on the patch-level alignment of our chosen diffusion features in Fig. 4, we assume that the normalized cut algorithm would provide sharp segments, each corresponding to a precise semantic concept as distinct objects would manifest as weakly connected components in a patch similarity matrix. Following this intuition, we construct an affinity matrix $\bm{W}$, by computing the cosine similarity between each pair of patches.  In addition, since NCut criterion assesses both the overall dissimilarity among different segments and the similarity within each segment, we opt to emphasize inter-segments dissimilarity by exponentiating each element with a positive integer value $\alpha$:
@@ 3.3 High-Resolution Concept Assignment
    Concept Assignment. A naive approach to obtain segmentation maps at the original image resolution consists in performing a nearest-neighbor upsampling. Despite its straightforwardness, this approach results in a blocky output structure as all pixels within the same feature patch are assigned to the same concept. Alternatively, we opt to first bilinearly upsample our low-resolution feature map $\hat{z}$ to match the original image spatial size and then proceed with the pixel/concept assignment. Specifically, for each concept $c_{i}\in\mathcal{C}$, we compute its cosine similarity with the upsampled features $\hat{z}_{up}$. This yields a similarity matrix of size $(H\times W\times K)$ where $K=|\mathcal{C}|$. Then, the assignment process simply consists in taking the argmax across the $K$ channels. The obtained segmentation map $\mathcal{S}$ is eventually refined with a pixel-adaptive refinement module [69].
@@ 4.1 Results on Zero-shot Segmentation
Model LD AX UA VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K Extra-Training    IIC [19] 9.8 - - 6.7 6.4 -    MDC [76] - - - 9.8 7.1 -    PiCIE [21] - - - 13.8 12.3 -    PiCIE+H [21] - - - 14.4 -    STEGO [20] - - - 28.2 21.0 -    ACSeg [18] 53.9 - - 28.1 - - Training-free    ReCO [53] 25.1 19.9 15.7 26.3 19.3 11.2    MaskCLIP [52] 38.8 23.6 20.6 19.6 10.0 9.8    MaskCut $(k=5)$ [32] 53.8 43.4 30.1 41.7 18.7 35.7    DiffSeg [1] - - - 43.6 21.2 -    DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7    DiffCut (Ours) 65.2 56.5 34.1 49.1 30.6 44.3
With our set of default hyperparameters, DiffCut significantly outperforms all other baselines despite not relying on language dependency, auxiliary images or unsupervised adaptation. On average, our method achieves a gain of +7.3 mIoU over the second best baseline. Notably, DiffCut exceeds MaskCut with an average improvement of +9.4 mIoU. Additionally, it outperforms the previous state-of-the-art method in unsupervised segmentation, DiffSeg, by +5.5 mIoU on COCO-Stuff and +9.4 mIoU on Cityscapes. The superiority of DiffCut over these two methods demonstrates our two key contributions: the high quality of our visual features for semantic segmentation and the flexibility of the recursive NCut algorithm in adjusting the number of segments according to the visual content of each image. The effectiveness of our method is further corroborated by our qualitative results shown in Fig. 1. In comparison to DiffSeg, DiffCut provides finely detailed segmentation maps that more closely align with semantic concepts. Additional examples can be found in Supplementary L.
We note here that, as the granularity of annotations varies across target datasets, our fixed set of hyperparameters can not be in the optimal regime on each of them. Therefore, relaxing the condition on prior knowledge about the target dataset, we report in Supplementary G results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images from the target training split.
@@ 4.2 Semantic Coherence in Vision Encoders
As in [79], we collect patch representations from various vision encoders and store their corresponding target classes using the segmentation labels. Given $\hat{z}_{i}=\mathcal{E}(\mathbf{x_{1}})_{i}\in\mathbb{R}^{D_{v}}$ and $\hat{z}_{j}=\mathcal{E}(\mathbf{x_{2}})_{j}\in\mathbb{R}^{D_{v}}$, the patch representations of images $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ at respectively index $i$ and $j$, we compute their cosine similarity  and use this score as a binary classifier to predict if the two patches belong to the same class. Given $l(\mathbf{x_{1}})_{i}$ and $l(\mathbf{x_{2}})_{j}$, the labels associated to the patches, if $l(\mathbf{x_{1}})_{i,j}=l(\mathbf{x_{2}})_{p,q}$, the target value for binary classification is 1, else 0. We present in Fig. 3 the ROC curve and AUC score for our candidate models. We observe that SSD-1B UNet encoder [35] demonstrates a greater patch-level alignment than any other candidate model with an AUC score of 0.83, even surpassing DINOv2 [5]. We further exhibit the outstanding alignment between patch representations associated to semantically similar concepts with qualitative results in Fig. 4. We provide additional qualitative examples patch-level alignment in Supplementary K.
@@ 4.3 Ablation study
Table 2: Ablation Study. The recursive partitioning of DiffCut yields superior results to both the self-attention merging process of DiffSeg and Automated Spectral Clustering.
@@
Recursive Normalized Cut vs Automated Spectral Clustering.
@@ 4.4 Model Analysis
Figure 5: Sensitivity of DiffCut. As $\alpha$ increases, DiffCut shows competitive results for a broad range of $\tau$ values.
Hyperparameters Impact.
@@
Model LD VOC Context COCO-Object Extra-Training    ViL-Seg [82] 37.3 18.9 -    TCL [50] 55.0 30.4 31.6    CLIPpy [48] 52.2 - 32.0    GroupVIT [49] 52.3 22.4 24.3    ViewCo [51] 52.4 23.0 23.5    SegCLIP [54] 52.6 24.7 26.5    OVSegmentor [83] 53.8 20.4 25.1 Training-free    ReCO [53] 25.1 19.9 15.7    MaskCLIP [52] 38.8 23.6 20.6    CLIP-DIY [84] 59.9 19.7 31.0    FreeSeg-Diff [85] 53.3 - 31.0    DiffCut 63.0 24.6 36.0
@@ Appendix A Recursive Normalized Cut on Diffusion Features
  • Encode the image $\mathcal{I}$ with the VAE of the diffusion model: $z=\mathcal{E}_{\text{VAE}}(\mathcal{I})$
@@ Appendix B Impact of PAMR
Appendix C Additional Comparison with MaskCut
DiffCut, unlike MaskCut [32], is capable of providing dense segmentation maps and dynamically adapting the number of detected segments based on the visual content of an image. In contrast, MaskCut can only detect a fixed number of segments, making it less suitable for image segmentation. This limitation arises from MaskCut’s use of an iterative graph partitioning approach, where graph nodes associated with detected objects are masked. As a result, each segment is treated as a single object and cannot be refined after detection, which severely restricts its ability to identify a large number of objects. To highlight the superiority of our recursive partitioning over MaskCut’s iterative process, we present below a comparison between DiffCut and MaskCut with $k$ the number of objects to be detected varying in $\{3,5,20\}$.
Table 6: Comparison with MaskCut. DiffCut recursive partitioning algorithm yields superior results than MaskCut iterative partitioning.
Model VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K DiffCut 62.0 54.1 32.0 46.1 28.4 42.4 MaskCut $(k=3)$ 53.7 42.3 30.9 41.8 18.0 33.7 MaskCut $(k=5)$ 53.8 43.4 30.1 41.7 18.7 35.7 MaskCut $(k=20)$ 53.8 43.5 30.0 41.5 18.0 35.6
DiffCut significantly outperforms MaskCut, regardless of the chosen value of $k$. DiffCut’s improvement showcases our two key contributions: the effectiveness of our visual features for semantic segmentation and the ability of the recursive NCut algorithm to dynamically adjust the number of segments based on each image’s visual content.
Appendix D DiffCut with Alternative Diffusion Backbones
To further display the relevance of diffusion features, we show that DiffCut achieves competitive performance even when using smaller diffusion backbones than SSD-1B. Specifically, we test two alternatives: SD1.4 and SSD-Vega [35] (another distilled version of SDXL). The UNet encoder in SD1.4 has 260M parameters, comprising approximately 30% of the overall UNet, while the UNet encoder in SSD-vega has 240M parameters, making up around 32% of the UNet.
Table 7: Performance of DiffCut with alternative diffusion backbones.
Model VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K SD1.4 57.5 52.8 30.0 45.2 24.5 36.7 SSD-Vega 62.2 56.4 34.9 49.5 30.1 45.7 SSD-1B 65.2 56.5 34.1 49.1 30.6 44.3 DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7
The results obtained using these two backbones are consistent with those achieved with SSD-1B. While the SD1.4 UNet encoder shows a slight performance drop compared to SSD-1B, DiffCut still significantly outperforms DiffSeg. Notably, with the SSD-Vega UNet encoder, DiffCut delivers performance comparable to SSD-1B, despite having only half the number of parameters.
Appendix E Mask Upsampling
@@
Appendix F Visual Encoders KMeans Comparison
@@
Appendix G Hyperparameter $\tau$ tuning
@@
Appendix H Image Noising
@@
Appendix I Hyperparameter Sensitivity
@@
Appendix J Visualization of the effect of $\tau$
@@
Appendix K Semantic Coherence in Vision Encoders
@@
Appendix L Additional Visualization
@@
Appendix M Datasets Licenses